# Calibration of Deep Learning Classification Models in fNIRS

### 23 Feb 2024 v1

@@ I INTRODUCTION
There are various methods for humans to understand brain signals, including electroencephalogram (EEG) [1], functional magnetic resonance imaging (fMRI) [2], and functional near-infrared spectroscopy (fNIRS) [3]. Functional Near-Infrared Spectroscopy (fNIRS) has become valuable for non-invasive monitoring of brain activity owing to its portability and reduced susceptibility to electrical noise and motion artifacts [4]. By detecting changes in oxygenated hemoglobin (HbO) and deoxygenated hemoglobin (HbR) concentrations via near-infrared light absorption, fNIRS facilitates the examination of human brain activity [3]. Understanding of fNIRS signals, particularly the classification of human behavioral intentions associated with these signals, holds promise for the advancement of brain-computer interfaces (BCIs) [5] and providing assistance for individuals with disabilities [6].
Currently, various effective methods exist for classifying fNIRS signals, including traditional machine learning and deep learning. Shin et al. [7] introduced Linear Discriminant Analysis (LDA), mapping fNIRS signals to distinct brain functional states. Chen et al. [8] proposed a feature extraction method based on the general linear model (GLM) to identify features from sparse fNIRS channels. However, traditional machine learning methods heavily rely on manual feature selection and prior knowledge, often resulting in poorer generalization compared with deep learning. Consequently, there is a growing trend in leveraging robust deep learning methods for fNIRS classification. Lyu et al. [9] applied CNN-based and LSTM-based methods to extract fNIRS signal characteristics. Furthermore, a 1D-CNN neural network [10] structure was employed to extract features from NIRS signals, enhancing classification accuracy by Sun et al. Rojas et al. [11] utilized a Bi-LSTM model to automatically extract features from raw fNIRS data. What’s more, Wang et al. [12] proposed a transformer-based classification network (fNIRS-T), to enhances data utilization and network representation capabilities by exploring spatial-level and channel-level signal representations for better classification. Additionally, Wang et al. [13] proposed incorporating delayed hemodynamic response as domain knowledge into fNIRS classification, presenting fNIRSNet as a straightforward yet effective model to enhance classification accuracy.

### 20 Mar 2024 v2

@@ I INTRODUCTION
There are various methods for humans to understand brain signals, including electroencephalogram (EEG) [1], functional magnetic resonance imaging (fMRI) [2], and functional near-infrared spectroscopy (fNIRS) [3]. Functional Near-Infrared Spectroscopy (fNIRS) has become valuable for non-invasive monitoring of brain activity owing to its portability and reduced susceptibility to electrical noise and motion artifacts [4]. By detecting changes in oxygenated hemoglobin (HbO) and deoxygenated hemoglobin (HbR) concentrations via near-infrared light absorption, fNIRS facilitates the examination of human brain activity [3]. Understanding of fNIRS signals, particularly the classification of human behavioral intentions associated with these signals, holds promise for the advancement of brain-computer interfaces (BCIs) and providing assistance for individuals with disabilities [5].
Currently, various effective methods exist for classifying fNIRS signals, including traditional machine learning and deep learning. Shin et al. [6] introduced Linear Discriminant Analysis (LDA), mapping fNIRS signals to distinct brain functional states. Chen et al. [7] proposed a feature extraction method based on the general linear model (GLM) to identify features from sparse fNIRS channels. However, traditional machine learning methods heavily rely on manual feature selection and prior knowledge, often resulting in poorer generalization compared with deep learning. Consequently, there is a growing trend in leveraging robust deep learning methods for fNIRS classification. Lyu et al. [8] applied CNN-based and LSTM-based methods to extract fNIRS signal characteristics. Furthermore, a 1D-CNN neural network [9] structure was employed to extract features from NIRS signals, enhancing classification accuracy by Sun et al. Rojas et al. [10] utilized a Bi-LSTM model to automatically extract features from raw fNIRS data. What’s more, Wang et al. [11] proposed a transformer-based classification network (fNIRS-T), to enhances network representation capabilities by exploring spatial-level and channel-level signal representations. Cao et al. [12] proposed integrating metric learning into fNIRS research to improve networks capability in excluding out-of-distribution outliers. Additionally, Wang et al. [13] proposed incorporating delayed hemodynamic response as domain knowledge into fNIRS classification to enhance classification accuracy.